OpenStudy (anonymous):

Find area of shaded section which is made by cutting a small semicircle out of a large semicircle. Assume the large semicircle is twice the radius of the smaller one. Round to nearest tenth...

OpenStudy (anonymous):



OpenStudy (campbell_st):

oops forgot the half \[A = 1/2 \pi \times 12^2 - 1/2 \pi \times 6^2\]